OpenStudy (anonymous):

3 • [6 • (8 – 12 ÷ 2)] + 5 • 6

OpenStudy (anonymous):

Ok so let's use order of operations First is parentheses, what do you get when you do what is inside the parentheses

OpenStudy (anonymous):

12 divided by 2=6 -8

OpenStudy (anonymous):

No look at it this way (8-12/2) You said 12/2=6 So it becomes (8-6) right?

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

So what is 8-6?

OpenStudy (anonymous):

2

OpenStudy (anonymous):

So now the entire thing is \[3\times(6\times2)+5\times2\]

OpenStudy (anonymous):

So let's do parentheses again What is 6 times 2?

OpenStudy (anonymous):

12

OpenStudy (anonymous):

So now we have \[3\times12+5\times6\]

OpenStudy (anonymous):

Do the two multiplication first Then add the results And there is your answer

OpenStudy (anonymous):

so the answer is 66

OpenStudy (anonymous):

Yup :)

OpenStudy (anonymous):

ok thank you